OpenStudy (anonymous):

please help: a bomb is dropped from an airplane 1,500m above the ground when the plane is moving horizontally at the rate of 160 km/hr.wher should the plane be with respect to the target when the bomb is dropped if a hit is to be made?

OpenStudy (amistre64):

if we neglect for air resistance and all the other stuff, except for gravity, that makes it more realistic ... the forward momentum stays constant; so if we do not put in a time constraint, the thing just continues to move forward until something blocks its path. Gravity is the only thing that i will be concentrating on at the moment, since that is the downward pull that makes it eventually hit the earth in a given time frame. And since there is no need to reinvent the wheel; lets use a gravity constant that equates to meters: -4.9 t^2 (m/s) Since the thing is dropped and not pushed, the initial velocity in the downward direction is 0. -4.9 t^2 +0 t And we start at an initial height of 1500 meters. -4.9 t^2 +1500; when we equate this a value of zero; we have hit the ground. It is a position function. The time it takes to hit the ground is then: 1500 = 4.9 t^2 1500/4.9 = t^2 sqrt(1500/4.9) = t =abt 17.5 seconds. --------------------------------------- convert km/h to m/s 160 km 1000 m 1 hour ------- * -------- * ------ = 2666' 2/3 (m/s) hour 1 km 60 sec ------------------------------------------- 2666' 2/3 * 17.5 = horisontal distance the thing will travel till it hits the ground.

OpenStudy (amistre64):